OpenStudy (anonymous):

Factor the following expression. 198q^3r^2 – 184q^2r^2 + 18qr^2

OpenStudy (anonymous):

Alright. What do you think the GCF is here?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

q and r^2

OpenStudy (anonymous):

Very good! :D \[2qr^{2}(\frac{198q^{3}r^{2}}{2qr^{2}} - \frac{184q^{2}r^{2}}{2qr^{2}} + \frac{18qr^{2}}{2qr^{2}}) = ?\]

OpenStudy (anonymous):

now what do I do?

OpenStudy (anonymous):

Divide the fractions.

OpenStudy (anonymous):

how do I divide the letters?

OpenStudy (anonymous):

Dividing the letters. Here's a hint: \[\huge\frac{a^{x}}{a^{y}} = a^{x - y}\]

OpenStudy (anonymous):

2qr^2(99q^2 - 92q+9)

OpenStudy (anonymous):

YAY :D 

OpenStudy (anonymous):

no what?

OpenStudy (anonymous):

no what do I do?

OpenStudy (anonymous):

i meant now

OpenStudy (anonymous):

is that the final answer?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thanks